I have tryed and tryed and I cant get it! Its -40-52=?

Found 3 solutions by Earlsdon, JimMarshall, Alan3354:
Answer by Earlsdon(6294)

You probably need to review operations with signed numbers.
-40 - 52
Remember the rule: "To subtract an integer, you add its additive inverse."
So, -40 - 52 = -40 + (-52) Adding integers with the same sign (here, they're both negative) just add their absolute values and the answer gets the same sign as the integers.
-40 - 52 = -40 + (-52) = -92

Answer by JimMarshall(82)

-40 - 52 = -40 + (-52) = -92
Ans

Answer by Alan3354(69443)

-40-52=?
----------
-1*40 + -1*52
= -1*(40 + 52)
= -1*92
= -92